OpenStudy (anonymous):

Suppose that $823.25 is invested in a savings account with an APR of 12% compounded monthly. What is the future value of the account in 5 years?

OpenStudy (anonymous):

do you know the formula for this?

OpenStudy (anonymous):

Yes I dont think im setting it up correctly. Because it is a multiple choice question

OpenStudy (anonymous):

should look like \[823.25\times (1+\frac{.12}{12})^{12\times 5}\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

or \[823.25(1.01)^{60}\]

OpenStudy (anonymous):

yw